### v2

@@ Abstract
benchmark, neuromorphic computing, computational neuroscience, energy consumption, performance optimization, cortical microcircuit, spiking neural network model
@@
Neuroscience increasingly depends on mathematical modeling and simulations to bridge the gap between the properties of single neurons and the function of the entire brain [2]. Thereby, neuroscience moves towards digital twins integrating anatomical and physiological data [3, 4, 5] and enabling virtual experiments that cannot be done in organisms. A particular structure of interest in this process is the so called canonical microcircuit [6, 7, 8] in the neocortex. The term refers to the local cortical network, the evolutionary youngest part in the mammalian brain [9]. The microcircuit is canonical in the sense that its basic architecture is conserved across species and modalities [10], making it a natural target for the search of generic computational principles in the brain. In 2014 a model was proposed representing all the nerve cells (neurons) and all their contact points (synapses) below the surface area of of cortex [11]; that is a network size of on the order of one hundred thousand neurons and one billion synapses. This removed all uncertainties about the effects of downscaling on network activity in earlier models [12]. The model represents four cortical layers by populations of excitatory and inhibitory leaky integrate-and-fire model neurons (network sketch in Fig. 1). Due to the simplicity of its components and the limited data available at the time, the explanatory scope of the model is limited, even though it reproduces basic features of cortical activity such as asynchronous and irregular spiking at biologically plausible population-specific rates. Initially the simulation results were confirmed by independent researchers using different simulation codes [13, 14]. The model and variations thereof have been used for various neuroscientific investigations (e.g., [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]) and as a test bed for theoretical methods (e.g., [28, 29, 30, 31, 32]). Nevertheless, when it was first proposed, run times of several minutes were required for the simulation of of biological time, limiting the practical use of the model. The opportunity to overcome a barrier in terms of network size, and the availability of a reproducible model combined with the computational challenge of bringing down simulation time of a microscopically parallel problem sparked a race in the computational neuroscience and neuromorphic computing communities for ever faster and more energy-efficient simulation. In particular the frequent updates (at least every of model-time in time-driven simulations) as well as the communication and delivery of a large number of events (also called spikes) during state propagation call for innovative solutions in hardware and software, beyond conventional approaches in High-Performance Computing (HPC). A few years later, the model has been simulated on a range of platforms including a large-scale neuromorphic system [1, 33], many-core CPU systems [1, 34] GPUs [35, 36, 37, 38] and FPGAs [39, 40]. In this time span, real-time performance was reached and simulation time and energy per synaptic event dropped by two orders of magnitude. The circuit can now be simulated an order of magnitude faster than real time. It is therefore time to gather the results distributed in the literature, contrast the findings and identify the advantages and bottlenecks of different approaches. From this we derive limitations of present technology and requirements for future neuromorphic architectures. From the perspective of computational science we discuss what made the PD14 model successful in becoming a de facto standard benchmark for our community and what obstacles the authors faced using the benchmark. Finally, we discuss next challenges for the simulation of large-scale neuronal networks and characterize desired properties of next generation benchmark models.
Results
@@
The PD14 has become a de facto standard benchmark. In the beginning, an executable model description was only available in the original simulation language (SLI) of the NEST code [43] and this has been included in all versions of NEST since v2.4.0 (released in 2015). But, thanks to progress in the community and large-scale projects, executable descriptions abstracted from a particular simulation engine became available. Nevertheless, a performance result is only meaningful if a benchmark reaches the same accuracy of the solution as the reference data. This is because generally, a lower accuracy can be reached with less computational effort. The neuroscientific purpose of the PD14 model is to reproduce certain characteristics of the activity of the respective biological neuronal network. Therefore, van Albada et al. [1] define a set of measures a benchmark of PD14 needs to fulfill. A later study [44] finds that the specified correlation measure is good enough to expose corrupted activity, but that model-specific correlation structure is only exhibited at considerably longer simulation times. This study also shows that the inhomogeneity of the network structure hides the distribution of synaptic weights: the correlation measures remain identical if all weights are collapsed to the mean value of the corresponding normal distribution. Retrospectively this is apparent as a mean-field theory [28] can quantitatively reproduce the first and second-order statistics of the population activity. Thus, the characterization of the solution does not require the simulation of a spiking neuronal network with an individual representation of all neurons and synapses. The synaptic weights can be replaced by a single value per neuron population [44] and connections can be generated on-the-fly with a generator for pseudo-random numbers [39]. This dramatically reduces the amount of memory a simulation engine needs to access. Nevertheless, so far only direct simulation delivers the full distributions of the statistical measures. It remains to the good will of the scientists to carry out the benchmark with simulation engines that are in principle capable of representing synaptic weights as dynamical variables as required for plasticity and learning. Spike-based plasticity rules directly interact with the correlation structure. If constraints of the engine at hand do not allow a representation at this level of resolution, the results may still be interesting but the conditions need to be declared.
@@ Discussion
Unlike other benchmarks in the area of brain-inspired computing, our metrics for comparing the performance of different approaches are not addressing the accuracy achieved in solving a particular task. It is already part of the verification of a certain system that a minimal accuracy is met. The actual comparison is on the time it takes to reach the solution (at required accuracy) and the energy consumed.
@@
Since the PD14 model’s conception, new data sets have emerged, expanding our neuroscientific understanding of cortical microcircuits. While model revisions could incorporate additional insights [4, 26, 59], they are not discussed here as they would not significantly alter the computational load. Still we argue for a transparent and systematic tracking of model revisions as done with the PD14 model within the NEST code over the past ten years and from now on in a self-contained Model reference repository.
@@
Slow processes like long-term plasticity, system-level learning, and development increase the computational costs of any model. Thus, at the scale of PD14 we will face longer simulation times than predicted by the data shown here, and plastic processes still cannot be studied in larger models even with state-of-the-art simulation engines. Another challenge therefore consists in devising and maturing additional representative benchmark models which require different metrics for evaluating the accuracy. It is here where benchmarks concentrating on time to solution and energy to solution converge with benchmarks concentrating on the performance in fulfilling a function [64].
Much like the supercomputing community has learned that a useful benchmarking suite needs to be multi-disciplinary for a multi-purpose system (see for example [53]), the field of neuromorphic computing needs to map out its application domain. The optimal application area of a given approach can then be characterized as a particular territory on this map.
@@
Methods
Technology and study details
Here we elaborate on the different technologies and respective studies contributing to the race. The focus is on facts that differentiate performance results obtained on different systems, assuming that simulations achieve a sufficient accuracy. For the complete study descriptions, we refer to the original publications.
Two different sets of Initial conditions have been employed in the simulations resulting in statistically identical network activity but differing in the transient activity: the original initial conditions used in [11] and amended initial conditions suggested in [33].
NEST CPU
NEST [43] is a simulation software for spiking neural networks exploiting conventional CPU-based high-performance computing infrastructure. NEST is written in C++, uses OpenMP [65] and MPI [66] for parallelization, and has a Python user interface PyNEST [67, 68] (source code: https://github.com/nest/nest-simulator). The code represents numerical values in double-precision floating-point format, and solves sub-threshold dynamics with exact integration [69]. The network construction and simulation phase are natively executed in direct succession and both parallelized across all resources of the same system. In the studies considered here, power of NEST simulations is measured with rack Power Distribution Units (PDUs), remotely reading the active power of the compute notes under load approximately once per second.
@@ Methods
GeNN
GPU enhanced Neuronal Networks (GeNN) [45] is a C++ library for generating efficient GPU kernels for spiking neural network simulation. In more recent versions, this library is exposed to Python as PyGeNN [37]. GeNN is designed to be highly flexible and one of the key enablers of this flexibility is that all neuron and synapse models are specified in a C-like language (GeNNCode) which can be written directly in Python model descriptions. While GeNN supports data-parallel training of spiking machine learning models across multiple GPUs using NCCL [75], its current focus is performance on single GPUs.
@@
This study presents extensions to GeNN that support heterogeneous dendritic delays and on-GPU initialisation of state variables and connectivity (specified in GeNNCode). The authors benchmark a C++ implementation [76] of the PD14 model on GeNN 3.2.0 [77] and a variety of GPU hardware (Jetson TX2, GeForce GTX 1050Ti, Tesla K40c and Tesla V100). The simulations use the original initial conditions, record all spikes, and the external drive consists of Poisson spike trains, directly delivered to each neuron. The benchmarks cover the duration of and separately assess the wall clock time required for initialization and state propagation. The authors record the power usage over time at the main socket and report energy to solution as well as energy per synaptic event. The fastest simulation is achieved on the V100 (KN18). The investigation did not have physical access to the V100 system and therefore estimates power consumption by subtracting the GPU power reported by nvidia-smi on the K40-based system and adding on the values for the V100.
@@
SpiNNaker
SpiNNaker [46] is a neuromorphic hardware platform designed to run spiking neural network simulations in biological real-time. This highly-parallel architecture uses low-power ARM-968 cores running at 200 Mhz, with 64KB local data memory and 32 KB local instruction memory, and 32-bit fixed-point calculations. Thus, there is no hardware support for floating-point operations. SpiNNaker chips made of up to 18 such cores are organized into boards of 48-chips which can be connected together. The biggest SpiNNaker machine built to date consists of over 1-million cores operating as a single machine. SpiNNaker machines execute spiking neural network simulations using the PyNN interface and sPyNNaker [80], and SpiNNTools [81] software APIs for facilitating problem description, partitioning, mapping, execution, and results extraction. To measure energy use in the studies considered here, a wall-socket power meter monitors an entire SpiNNaker system: SpiNNaker boards, communications switch, power supplies, and cooling fans. The meter used to profile PD14 simulations provides a reading accurate to 0.01 kWh, with measurements taken at the beginning and end of simulations via software controlled readings. The authors subsequently convert this energy to solution into energy per synaptic event by dividing by the total number of synaptic events, and assuming constant power consumption throughout the simulation period [1, 33].
@@
This study completely expresses the system architecture of CsNN in a high-level programming language and maps the specification to FPGA logic using high-level synthesis tools. This design flow enables the implementation and characterization of a large number of architecture variants as part of a broad design space exploration. The authors employ the framework to characterize different advanced ODE solvers in terms of their performance, with the perspective of integrating more realistic neuron and synapse models as used in PD14. Non-linear equations arise, for example, when considering bio-realistic conductance-based synaptic couplings, AdEx neurons [85], Izhikevich neurons [86], or a non-linear membrane dynamics based on ion channels. In particular, all arithmetic operators are implemented in single-float precision. The work also extends the concept of procedural connectivity [87] to the representation of synaptic connections with the attributes of synaptic weight and delay. Instead of storing synaptic parameters in a memory system, the algorithm derives all connection parameters at runtime by addressing individual seeds. The authors propose dedicated random number generators for the sampling from arbitrary distribution functions for the runtime-efficient online generation of synaptic parameters. The approach maps the connectivity of PD14 to a highly parallel near-memory computing architecture with a high compression factor and rapid look-up of individual connections.
@@
Model reference
The original neuroscientific publication introducing the PD14 model and describing the data integration process as well as the neuroscientific conclusions is [11]. However, on the basis of the progress of the community in the formal description of neuronal network models since the first publication of PD14 [2, 88, 94] and the results of the present study, we here recommend to use the novel model reference: The open access web site https://github.com/INM-6/microcircuit-PD14-model provides a detailed, implementation-agnostic mathematical description of the model, as well as a documented PyNEST reference implementation.
Model and simulation parameters

### v3

@@ Abstract
benchmark, neuromorphic computing, computational neuroscience, energy consumption, performance optimization, cortical microcircuit, spiking neural network model1 Introduction
@@
Neuroscience increasingly depends on mathematical modeling and simulations to bridge the gap between the properties of single neurons and the function of the entire brain [2]. Thereby, neuroscience moves towards digital twins integrating anatomical and physiological data (e.g., using rat [3, 4] or mouse [5] models) and enabling virtual experiments that cannot be done in organisms. A particular structure of interest in this process is the so called canonical microcircuit [6, 7, 8] in the neocortex. The term refers to the local cortical network, the evolutionary youngest part in the mammalian brain [9]. The microcircuit is canonical in the sense that its basic architecture is conserved across species and modalities [10], making it a natural target for the search of generic computational principles in the brain. In 2014 a model was proposed representing all the nerve cells (neurons) and all their contact points (synapses) below the surface area of of cortex [11]; that is a network size of on the order of one hundred thousand neurons and one billion synapses. This removed all uncertainties about the effects of downscaling on network activity in earlier models [12]. Downscaling here refers to the concept of reducing the number of neurons and synapses in a model for computational tractability, while attempting to preserve the dynamics and keeping other structural characteristics like the probability of two neurons forming a connection invariant. It does not refer to representing neurons and synapses at a finer level of description, like the level of electrical compartments or the molecular level. The model represents four cortical layers by populations of excitatory and inhibitory leaky integrate-and-fire model neurons (network sketch in Fig. 1). Due to the simplicity of its components and the limited data available at the time, the explanatory scope of the model is limited, even though it reproduces basic features of cortical activity such as asynchronous and irregular spiking at biologically plausible population-specific rates. Initially the simulation results were confirmed by independent researchers using different simulation codes [13, 14]. The model and variations thereof have been used for various neuroscientific investigations (e.g., attention [15, 16], orientation selectivity [17], inhibitory neuron types [18, 19, 20], network connectivity [21, 22], building block of multi-area models [23, 24], forward-modeling of extracellular potentials [25, 26]) and as a test bed for theoretical methods (e.g., mean-field and linear response theory [27, 28], population density models [29, 30], stochastic population equations [31, 32]). Nevertheless, when it was first proposed, run times of several minutes were required for the simulation of of biological time, limiting the practical use of the model. A recent review [33] analyses the impact of the model on the fields of computational neuroscience and neuromorphic computing.
The opportunity to overcome a barrier in terms of network size, and the availability of a reproducible model combined with the computational challenge of bringing down simulation time of a microscopically parallel problem sparked a concrete and quantifiable race in the computational neuroscience and neuromorphic computing communities for ever faster and more energy-efficient simulation. In particular the frequent updates (at least every of model-time in time-driven simulations) as well as the communication and delivery of a large number of events (also called spikes) during state propagation call for innovative solutions in hardware and software, beyond conventional approaches in High-Performance Computing (HPC). A few years later, the model has been simulated on a range of platforms including a large-scale neuromorphic system (SpiNNaker [1, 34]), many-core CPU systems (NEST [1, 35]) GPUs (GeNN [36, 37], NEST GPU [38, 39] and FPGAs (CsNN [40], neuroAIx [41]). In this time span, real-time performance was reached and simulation time and energy per synaptic event dropped by two orders of magnitude. The circuit can now be simulated an order of magnitude faster than real time. The role of the model as a benchmark was never declared by any organization but emerged over the years. The studies reproducing the work were done independently by different laboratories. It is therefore time to gather the results distributed in the literature, contrast the findings and identify the advantages and bottlenecks of different approaches. From this we derive limitations of present technology and requirements for future neuromorphic architectures. From the perspective of computational science we discuss what made the PD14 model successful in becoming a de facto standard benchmark for our community and what obstacles the authors faced using the benchmark. Finally, we discuss next challenges for the simulation of large-scale neuronal networks and characterize desired properties of next generation benchmark models.
2 Results
@@
The PD14 has become a de facto standard benchmark. In the beginning, an executable model description was only available in the original simulation language (SLI) of the NEST code [44] and this has been included in all versions of NEST since v2.4.0 (released in 2015). But, thanks to progress in the community and large-scale projects, executable descriptions abstracted from a particular simulation engine became available. Nevertheless, a performance result is only meaningful if a benchmark reaches the same accuracy of the solution as the reference data. This is because generally, a lower accuracy can be reached with less computational effort. The neuroscientific purpose of the PD14 model is to reproduce certain characteristics of the activity of the respective biological neuronal network. Therefore, [1] define a set of measures a benchmark of PD14 needs to fulfill. A later study [45] finds that the specified correlation measure is good enough to expose corrupted activity, but that model-specific correlation structure is only exhibited at considerably longer simulation times. This study also shows that the inhomogeneity of the network structure hides the distribution of synaptic weights: the correlation measures remain identical if all weights are collapsed to the mean value of the corresponding normal distribution. Retrospectively this is apparent as a mean-field theory [27] can quantitatively reproduce the first and second-order statistics of the population activity. Thus, the characterization of the solution does not require the simulation of a spiking neuronal network with an individual representation of all neurons and synapses. The synaptic weights can be replaced by a single value per neuron population [45] and connections can be generated on-the-fly with a generator for pseudo-random numbers [40]. This dramatically reduces the amount of memory a simulation engine needs to access. Nevertheless, so far only direct simulation delivers the full distributions of the statistical measures. It remains to the good will of the scientists to carry out the benchmark with simulation engines that are in principle capable of representing synaptic weights as dynamical variables as required for plasticity and learning. Spike-based plasticity rules directly interact with the correlation structure. If constraints of the engine at hand do not allow a representation at this level of resolution, the results may still be interesting but the conditions need to be declared.
@@ 3 Discussion
Unlike other benchmarks in the area of brain-inspired computing, our metrics for comparing the performance of different approaches are not addressing the accuracy achieved in solving a particular task. It is already part of the verification of a certain system and a precondition that a minimal accuracy is met. The actual comparison is on the time it takes to reach the solution (at required accuracy) and the energy consumed.
@@
Since the PD14 model’s conception, new data sets have emerged, expanding our neuroscientific understanding of cortical microcircuits. While model revisions could incorporate additional insights (e.g., target-type specific innervation preferences depending on whether a neuron is excitatory or inhibitory [60], distinct interneuron types and short-term plasticity [20], and point-neuron vs. multi-compartment neurons [5]), they are not discussed here as they would not significantly alter the computational load. Still we argue for a transparent and systematic tracking of model revisions as done with the PD14 model within the NEST code over the past ten years and from now on in a self-contained Model reference repository.
@@
More detailed neuron models and slow processes like long-term plasticity, system-level learning, and development increase the computational costs of any model. Thus, at the scale of PD14 we will face longer simulation times than predicted by the data shown here, and plastic processes still cannot be studied in larger models even with state-of-the-art simulation engines. Another challenge therefore consists in devising and maturing additional representative benchmark models which require different metrics for evaluating the accuracy. It is here where benchmarks concentrating on time to solution and energy to solution converge with benchmarks concentrating on the performance in fulfilling a function [65].
Much like the supercomputing community has learned that a useful benchmarking suite needs to be multi-disciplinary for a multi-purpose system (see for example [54]). At the same time not all benchmarks are of relevance for all systems. A chip for edge computing has a different area of application than a chip designed to simulate a cortical network. The field of neuromorphic computing needs to map out its application domain and arrive at a consistent set of benchmarks [65]. The optimal application area of a given approach can then be characterized as a particular territory on this map.
@@
4 Methods
4.1 Technology and study details
Here we elaborate on the different technologies and respective studies contributing to the race. We start with a general introduction to each system with main references, links to website and source code if available, and design goals. Regarding implementation languages we distinguish between the user interface and low-level components. The hardware is described as either the conventional target processor or a custom architecture; for an overview refer to Fig. 3. We also specify the parallelization model for software and hardware if applicable. A characterization of the employed spiking neural network simulation strategy includes the sequence of steps (e.g., code generation, compilation, network construction, and state propagation), but also details on algorithms and numerics. Each section concludes with details on power measurement. These brief simulator profiles are by no means complete as some approaches are the results of decades of development and user experience. Features only touched upon may be the support for certain neuron and synapse models or plasticity mechanisms, or scalability with respect to network size. Our focus is on aspects that differentiate PD14 performance results on different systems, assuming that simulations achieve a sufficient accuracy. Where applicable, we highlight what we have learned from a computational point of view by contributing to the race. Following the generic introductory paragraphs, we provide details on the particular PD14 results, i.e., the data points depicted in Fig. 2, complementary to the performance data and system specifications in Table 1. For the complete study descriptions, we refer to the original publications.
With the maturation of the model specification, the studies employ two different sets of Initial conditions resulting in statistically identical network activity but differing in the transient activity: the original initial conditions used in [11] and amended initial conditions suggested in [34].
4.1.1 NEST CPU
NEST [44] is a simulation software for spiking neural networks exploiting conventional CPU-based high-performance computing infrastructure. NEST is written in C++, uses OpenMP [66] and MPI [67] for parallelization, and has a Python user interface PyNEST [68, 69] (website: https://www.nest-simulator.org, source code: https://github.com/nest/nest-simulator). The code represents numerical values in double-precision floating-point format, and solves sub-threshold dynamics with exact integration [70]. The network construction and simulation phase are natively executed in direct succession and both parallelized across all resources of the same system. In the studies considered here, power of NEST simulations is measured with rack Power Distribution Units (PDUs), remotely reading the active power of the compute notes under load approximately once per second.
@@ 4.1.2 NEST GPU
4.1.3 GeNN
GPU enhanced Neuronal Networks (GeNN) [46] is a C++ library for generating efficient GPU kernels in CUDA or HIP for spiking neural network simulation (website: https://genn-team.github.io, source code: https://github.com/genn-team/genn). In more recent versions, this library is exposed to Python as PyGeNN [37]. GeNN is designed to be highly flexible, and one of the key enablers of this flexibility is that all neuron and synapse models are specified in a C-like language (GeNNCode), which can be written directly in Python model descriptions. While GeNN supports data-parallel training of spiking machine learning models across multiple GPUs using NCCL [76], its current focus is performance on single GPUs. Spiking neural network simulations in GeNN comprise a code generation stage, compilation of generated C++ and CUDA/HIP code, followed by network initialisation and simulation. Network initialisation can largely be offloaded onto the GPU. Users of GeNN define neuron dynamics in terms of an update procedure for a timestep, meaning that the model definition already encompasses both the underlying differential equations and the numerical solver.
@@ 4.1.3 GeNN
This study presents extensions to GeNN that support heterogeneous dendritic delays and on-GPU initialisation of state variables and connectivity (specified in GeNNCode). The authors benchmark a C++ implementation [77] of the PD14 model on GeNN 3.2.0 [78] and a variety of GPU hardware (Jetson TX2, GeForce GTX 1050Ti, Tesla K40c and Tesla V100). The simulations use the original initial conditions, record all spikes, and the external drive consists of Poisson spike trains, directly delivered to each neuron. The benchmarks cover the duration of and separately assess the wall clock time required for initialization and state propagation. The authors record the power usage over time at the main socket and report energy to solution as well as energy per synaptic event. The fastest simulation is achieved on the V100 (KN18). The investigation did not have physical access to the V100 system and therefore estimates power consumption by subtracting the GPU power reported by nvidia-smi on the K40-based system and adding the values for the V100.
@@
4.1.4 SpiNNaker
SpiNNaker [47] is a neuromorphic hardware platform designed to run spiking neural network simulations in biological real-time (source code: https://github.com/spinnakermanchester). This highly-parallel architecture uses low-power ARM-968 cores running at 200 Mhz, with 64KB local data memory and 32 KB local instruction memory, and 32-bit fixed-point calculations. Thus, there is no hardware support for floating-point operations. SpiNNaker chips made of up to 18 such cores are organized into boards of 48-chips which can be connected together. The biggest SpiNNaker machine built to date consists of over 1-million cores operating as a single machine. SpiNNaker machines execute spiking neural network simulations using the PyNN interface and sPyNNaker [81], and SpiNNTools [82] software APIs for facilitating problem description, partitioning, mapping, execution, and results extraction. To measure energy use in the studies considered here, a wall-socket power meter monitors an entire SpiNNaker system: SpiNNaker boards, communications switch, power supplies, and cooling fans. The meter used to profile PD14 simulations provides a reading accurate to 0.01 kWh, with measurements taken at the beginning and end of simulations via software controlled readings. The authors subsequently convert this energy to solution into energy per synaptic event by dividing by the total number of synaptic events, and assuming constant power consumption throughout the simulation period [1, 34].
@@ 4.1.5 CsNN
This study completely expresses the system architecture of CsNN in a high-level programming language and maps the specification to FPGA logic using high-level synthesis tools. This design flow enables the implementation and characterization of a large number of architecture variants as part of a broad design space exploration. The authors employ the framework to characterize different advanced ODE solvers in terms of their performance, with the perspective of integrating more realistic neuron and synapse models as used in PD14. Non-linear equations arise, for example, when considering bio-realistic conductance-based synaptic couplings, AdEx neurons [86], Izhikevich neurons [87], or a non-linear membrane dynamics based on ion channels. In particular, all arithmetic operators are implemented in single-float precision. The work also extends the concept of procedural connectivity [88, 89] to the representation of synaptic connections with the attributes of synaptic weight and delay. Instead of storing synaptic parameters in a memory system, the algorithm derives all connection parameters at runtime by addressing individual seeds. The authors propose dedicated random number generators for the sampling from arbitrary distribution functions for the runtime-efficient online generation of synaptic parameters. The approach maps the connectivity of PD14 to a highly parallel near-memory computing architecture with a high compression factor and rapid look-up of individual connections.
@@ 4.2 Benchmarking recipe
4.2.1 Model reference
The original neuroscientific publication introducing the PD14 model and describing the data integration process as well as the neuroscientific conclusions is [11]. However, on the basis of the progress of the community in the formal description of neuronal network models since the first publication of PD14 (including the scientific case for large-scale brain simulations [2], connectivity concepts [90], and the modeling language NESTML [96]) and the results of the present study, we recommend to use the novel model reference: The open access web site https://github.com/INM-6/microcircuit-PD14-model provides a detailed, implementation-agnostic mathematical description of the model, as well as a documented PyNEST reference implementation. The repository also collects the performance results of this study and supports the inclusion of new data points as they become available.
4.2.2 Model and simulation parameters